OpenStudy (anonymous):

Need Help!! (-7,9)(-5,-3)

jimthompson5910 (jim_thompson5910):

What do you want to do here?

OpenStudy (anonymous):

slope (for example y=5x-19)

jimthompson5910 (jim_thompson5910):

If you want to find the slope of the line through the two points, then m = (y2-y1)/(x2-x1) m = (-3-9)/(-5-(-7)) m = (-3-9)/(-5+7) m = (-12)/2 m = -6 So the slope of the line that passes through the two points is -6

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

First draw a quick sketch.

OpenStudy (anonymous):

can you write the equation?

OpenStudy (skullpatrol):